# Model-Based Learning for DOA Estimation with One-Bit Single-Snapshot Sparse Arrays

### v1

@@ Abstract
Index Terms:
DOA estimation, single-snapshot, one-bit quantization, sparse arrays, off-grid, machine learning, majorization minimization
@@ I Introduction
In antenna array systems, high-precision, high-sampling-rate analog-to-digital converters (ADCs) are costly and power-intensive. One-bit quantization provides a cost-effective solution, simplifying the sampling hardware and lowering the sampling rates by reducing the number of bits per sample to just one [17]. Consequently, one-bit signal processing has drawn significant research interest in many fields. In communications, a novel framework for computing the minimum mean square error (MMSE) channel estimator in one-bit quantized multi-input multi-output (MIMO) systems was developed in [18]. In [19], a two-stage detection method for massive MIMO with one-bit ADCs was proposed, which employs Bussgang-based linear receivers and a model-driven neural network, achieving robust performance at significantly reduced complexity. In [20], a novel one-bit MIMO precoding scheme was proposed using spatial Sigma-Delta modulation to effectively control quantization errors utilizing favorable conditions of a massive MIMO system. In the field of radar signal processing, target parameter estimation based on one-bit quantized measurements has emerged as a prominent research topic [21, 22, 23]. In [24], one-bit ADC radar parameter estimation was formulated as a cyclic multivariate weighted least squares problem. A dimension-reduced generalized approximate message passing approach was introduced in [25] for one-bit frequency-modulated continuous-wave (FMCW) radar that mitigates high-order harmonic interference. Multiple sparse recovery algorithms were extended in [26] to a one-bit quantization setting, resulting in one-bit sparse learning via iterative minimization (1bSLIM), likelihood-based estimation of sparse parameters (1bLIKES), one-bit sparse iterative covariance-based estimation (1bSPICE), and one-bit iterative adaptive approaches (1bIAA), to generate automotive radar range-Doppler imaging exploiting one-bit quantization.
@@ I-A Relevant Works
Subspace-based methods exploiting one-bit quantized data, along with performance analyses, were presented for both ULAs and SLAs [2, 27, 28]. It is shown in [2] that the identifiability of one-bit DOA estimation with SLAs matches unquantized measurements and a conservative Cramér-Rao bound (CRB) for quantized systems is derived. In [28], MUSIC-based DOA estimation leverages an approximate one-bit covariance matrix, derived as a scaled version of the unquantized covariance matrix. In the past decade, methods leveraging sparse reconstruction [29] and compressed sensing (CS) [30] have emerged to address one-bit DOA estimation and sparse sensing, leading to various estimators, such as binary iterative hard thresholding (BIHT) [31], robust one-bit Bayesian compressed sensing (R1BCS) [32]. The generalized sparse Bayesian learning (Gr-SBL) algorithm for both single- and multi-snapshot scenarios [33] can operate effectively with SLAs. However, all these methods only consider on-grid signals, as they determine signal DOAs by peak searching over a fixed discrete angular spectrum.
@@ I-B Contributions
The rest of the paper is organized as follows. Section II introduces the one-bit DOA estimation signal model and provides a mathematical formulation for both on-grid and off-grid cases. In Section III, we derive two algorithm frameworks SBRI and SBRI-X for one-bit DOA estimation, based on the Sparse Bayesian Reweighted Iterative approach, addressing both on-grid and off-grid DOA estimation. Section IV presents model-based neural networks inspired by algorithm unrolling. Section V provides numerical results for performance evaluation. Finally, Section VI concludes the paper.
@@ III Sparse Bayesian Learning Framework for One-Bit DOA Estimation
In this section, we introduce a sparse Bayesian reweighted iterative (SBRI) framework and its augmented variant for the estimation of on-grid and off-grid signal DOAs based on one-bit quantized array data.
@@ III-A One-Bit On-Grid DOA Estimation via SBRI Framework
where $\Phi\left(\cdot\right)$ denotes the CDF of the standard normal distribution, $\overline{y}_{m}$ is the $m$-th element in $\mathbf{y}$, $\mathbf{a}_{m}^{\mathrm{T}}$ is the $m$-th row vector in $\mathbf{\mathcal{A}}$. Since $\sigma>0$ and scaling $\mathbf{x}$ by a positive constant does not affect the values of $\overline{\bm{y}}$ or the DOA estimation results, letting $\hat{\mathbf{x}}=\frac{\sqrt{2}}{\sigma}\mathbf{x}$, (8) is reformulated as:
$p\left(\overline{\mathbf{y}}|\hat{\mathbf{x}}\right)=\!\prod_{m=1}^{M}$ $\Phi\left(\Re(\overline{y}_{m})\Re(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)\Phi\left(\Im(\overline{y}_{m})\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right).$ (9)
@@
where the parameter $\gamma>0$. As $\alpha$ approaches 00, $p\left(\hat{\mathbf{x}}\right)$ sharply peaks at $\hat{\mathbf{x}}=0$, enforcing sparsity on the estimated $\hat{\mathbf{x}}$. Notice that, if we choose the prior PDF as $p\left(\hat{\mathbf{x}}\right)=\frac{1}{\left|\hat{\mathbf{x}}\right|^{2}+\varepsilon}$, where $\varepsilon>0$ is a small constant, we will derive the 1bSLIM approach proposed in [26]. Using Bayes’ rule, the MAP estimator for on-grid model is given as:
$\hat{\mathbf{x}}^{\star}=\mathrm{arg}\min_{\hat{\mathbf{x}}}\bigg{\{}-\mathrm{ln}\,p\left(\overline{\mathbf{y}}|\hat{\mathbf{x}}\right)-\mathrm{ln}\,p\left(\hat{\mathbf{x}}\right)\bigg{\}}.$ (11)
@@
$\mathcal{L}_{\mathrm{ongrid}}$ $=\sum_{m=1}^{M}\Bigg{\{}-\mathrm{ln}\,\Phi\left(\Re\left(\overline{y}_{m}\right)\Re(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)$
$\ \ -\mathrm{ln}\,\Phi\left(\Im\left(\overline{y}_{m}\right)\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)+\sum_{i=1}^{N}\frac{\gamma\left|x_{i}\right|^{\alpha}}{\alpha}+\textrm{const}\Bigg{\}}.$ (12)
@@
$\sum_{m=1}^{M}\!\left\{\!-\mathrm{ln}\,\Phi\left(\Re\left(\overline{y}_{m}\right)\!\Re(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)\!-\!\mathrm{ln}\,\Phi\left(\Im\left(\overline{y}_{m}\right)\!\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)\!\right\}$
@@
$-\Re(v_{m}^{k})\Re(\overline{y}_{m})\Re(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\!-\!\Im(v_{m}^{k})\Im(\overline{y}_{m})\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})+c^{\prime}\bigg{.}\bigg{\}},$ (13)
@@
The third term in (12) is non-convex, but can be relaxed to a convex form via smooth approximation[48]:
@@
where $\eta>0$ is a small constant. A smaller value of $\eta$ provides a closer (15) approximation but may reduce the robustness of the algorithm. Typically, $\eta$ is set to $10^{-6}$.
Algorithm 1 SBRI for One-Bit On-Grid DOA.
2:Output: $\hat{\mathbf{x}}^{\star}$
4:for $k\leftarrow 1$ to $T_{\max}$ do
5:     Update $\hat{\mathbf{x}}^{k}$ using the formula in (17)
6:     Update $\gamma^{k}$ using the formula in (19)
7:     if $\frac{\|\hat{\mathbf{x}}^{(k+1)}-\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}{\|\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}\leq\epsilon_{0}$ or $k\geq T_{\max}$ then
8:         break
9:     else
10:         $\hat{\mathbf{x}}^{k+1}\leftarrow\hat{\mathbf{x}}^{k}$
11:         $\gamma^{k+1}\leftarrow\gamma^{k}$
12:     end if
13:end for
14:return $\hat{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k}$
@@
$\hat{\mathbf{x}}^{\star}=\mathrm{arg}\min_{\hat{\mathbf{x}}}\bigg{\{}\frac{1}{2}\left\|\mathbf{\mathcal{A}}\hat{\mathbf{x}}-\mathbf{v}^{k}\right\|_{2}^{2}+\frac{\gamma}{\alpha}\sum_{i=1}^{N}\left(\left|\hat{x}_{i}\right|^{2}+\eta\right)^{\frac{\alpha}{2}}+c^{\prime}\bigg{\}},$ (16)
where $\mathbf{v}^{k}=[v_{1}^{k},\cdots,v_{M}^{k}]^{\mathrm{T}}\in\mathbb{C}^{M\times 1}$. Note that (16) represents a classic regularized least squares problem, and $\gamma$ is a regularization parameter. Therefore, we can utilize the iteratively reweighted least squares method[49], using the estimates obtained from the $k$-th iteration step, to compute $\hat{\mathbf{x}}$ at the $(k+1)$-th step as:
$\hat{\mathbf{x}}^{k+1}$ $=\left[\mathcal{A}^{\mathrm{H}}\mathcal{A}+\gamma^{k}\bm{\Lambda}^{-1}(\hat{\mathbf{x}}^{k})\right]^{-1}\mathcal{A}^{\mathrm{H}}\mathbf{v}^{k},$ (17)
where

### v2

@@ I Introduction
In antenna array systems, high-precision, high-sampling-rate analog-to-digital converters (ADCs) are costly and power-intensive. One-bit quantization provides a cost-effective solution, simplifying the sampling hardware and lowering the sampling rates by reducing the number of bits per sample to just one [17]. Consequently, one-bit signal processing has drawn significant research interest in many fields. In communications, a novel framework for computing the minimum mean square error (MMSE) channel estimator in one-bit quantized multi-input multi-output (MIMO) systems was developed in [18]. In [19], a two-stage detection method for massive MIMO with one-bit ADCs was proposed, which employs Bussgang-based linear receivers and a model-driven neural network, achieving robust performance at significantly reduced complexity. In [20], a novel one-bit MIMO precoding scheme was proposed using spatial Sigma-Delta modulation to effectively control quantization errors utilizing favorable conditions of a massive MIMO system. In the field of radar signal processing, target parameter estimation based on one-bit quantized measurements has emerged as a prominent research topic [21, 22, 23]. A two-stage, tensor-decomposition-based channel-estimation framework for MIMO systems with movable antennas was proposed in [24]. Successive transmit-receive sweeps are arranged so that the resulting pilot measurements are stacked into a structured third-order tensor, on which canonical polyadic (CP) decomposition is applied to uniquely recover angle-of-arrival/departure (AoA/AoD) and complex path gains, thereby enabling accurate channel reconstruction for arbitrary antenna positions while reducing pilot overhead far below that of compressed sensing (CS) baselines. The channel state information (CSI) bottleneck in XL-MIMO-OFDM systems with dynamic metasurface antennas is alleviated by microstrip-sequencing the pilots into a fourth-order tensor that admits CP decomposition [25]. By exploiting Vandermonde factor matrices, a two-stage closed-form procedure is then used to uniquely estimate AoA/AoD, delays, and gains with training overhead proportional only to the number of multipaths, thereby reducing pilots and improving accuracy relative to MMSE and CS baselines. In [26], one-bit ADC radar parameter estimation was formulated as a cyclic multivariate weighted least squares problem. A dimension-reduced generalized approximate message passing approach was introduced in [27] for one-bit frequency-modulated continuous-wave (FMCW) radar that mitigates high-order harmonic interference. Multiple sparse recovery algorithms were extended in [28] to a one-bit quantization setting, resulting in one-bit sparse learning via iterative minimization (1bSLIM), likelihood-based estimation of sparse parameters (1bLIKES), one-bit sparse iterative covariance-based estimation (1bSPICE), and one-bit iterative adaptive approaches (1bIAA), to generate automotive radar range-Doppler imaging exploiting one-bit quantization.
@@ I-A Relevant Works
Subspace-based methods exploiting one-bit quantized data, along with performance analyses, were presented for both ULAs and SLAs [2, 29, 30]. It is shown in [2] that the identifiability of one-bit DOA estimation with SLAs matches unquantized measurements and a conservative Cramér-Rao bound (CRB) for quantized systems is derived. In [30], MUSIC-based DOA estimation leverages an approximate one-bit covariance matrix, derived as a scaled version of the unquantized covariance matrix. In the past decade, methods leveraging sparse reconstruction [31] and CS [32] have emerged to address one-bit DOA estimation and sparse sensing, leading to various estimators, such as binary iterative hard thresholding (BIHT) [33], robust one-bit Bayesian compressed sensing (R1BCS) [34]. The generalized sparse Bayesian learning (Gr-SBL) algorithm for both single- and multi-snapshot scenarios [35] can operate effectively with SLAs. However, all these methods only consider on-grid signals, as they determine signal DOAs by peak searching over a fixed discrete angular spectrum.
@@
A dynamic maximum likelihood estimator was proposed in [39] and applied to both hybrid analog-digital and fully digital structures, with precise DOA estimates obtainable using either a golden section search method or a gradient search approach. A tensor-based approach was introduced in [40] for efficiently estimating both channel state information and target parameters in massive MIMO integrated sensing and communication systems. By parameterizing the communication channel and utilizing a shared training pattern, it formulates the problem as a CP decomposition. The unified algorithm estimates key parameters and incorporates a segment-based training pattern to mitigate beam squint effects, resulting in improved accuracy, resolution, and reduced training overhead.
@@ I-B Contributions
The rest of the paper is organized as follows. Section II introduces the one-bit DOA estimation signal model and provides a mathematical formulation for both on-grid and off-grid cases. In Section III, we derive two algorithm frameworks, Sparse Bayesian Reweighted Iterative (SBRI) and SBRI-X, for one-bit DOA estimation, addressing both on-grid and off-grid DOA estimation problems. Section IV presents model-based neural networks inspired by algorithm unrolling. Section V provides numerical results for performance evaluation. Finally, Section VI concludes the paper.
@@ III Sparse Bayesian Learning Framework for One-Bit DOA Estimation
In this section, we introduce the SBRI framework and its augmented variant (SBRI-X) for the estimation of on-grid and off-grid signal DOAs based on one-bit quantized array data. The core SBRI loop reformulates one-bit MAP inference with a sparsity prior into a majorization-minimization (MM)-based iteratively reweighted least squares (IRLS) procedure. At each iteration, we first construct a surrogate for the quantized CDF likelihood, compute a pseudo-measurement that effectively compensates for the sign nonlinearity, and then solve a closed-form IRLS update for the spectral coefficients. The regularization weight is automatically adjusted based on the norm of the current estimate.
SBRI-X extends this framework by introducing noise offsets and replacing the CDF with a smooth sigmoid function. The sigmoid term is majorized using a second-order (quadratic) upper bound, enabling an alternating optimization scheme that performs two simple updates: an efficient IRLS step for the spectral coefficients and a closed-form update for the noise offsets.
@@ III-A One-Bit On-Grid DOA Estimation via SBRI Framework
where $\Phi\left(\cdot\right)$ denotes the CDF of the standard normal distribution, $\overline{y}_{m}$ is the $m$-th element in $\overline{\mathbf{y}}$, $\mathbf{a}_{m}^{\mathrm{T}}$ is the $m$-th row vector in $\mathbf{\mathcal{A}}$. Since $\sigma>0$ and scaling $\mathbf{x}$ by a positive constant does not affect the values of $\overline{\bm{y}}$ or the DOA estimation results, letting $\hat{\mathbf{x}}=\frac{\sqrt{2}}{\sigma}\mathbf{x}$, (8) is reformulated as:
[…]
@@
$p\left(\hat{\mathbf{x}}\right)=\prod_{i=1}^{N}\mathrm{exp}\left(-\frac{\gamma\left|\hat{x}_{i}\right|^{\alpha}}{\alpha}\right),\ \ 0<\alpha\leq 1,$ (10)
@@
[…]
@@
[…]
$\ \ -\mathrm{ln}\,\Phi\left(\Im\left(\overline{y}_{m}\right)\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\right)\left.+\sum_{i=1}^{N}\frac{\gamma\left|x_{i}\right|^{\alpha}}{\alpha}+\textrm{const}\right\}.$ (12)
@@
[…]
@@
$-\Re(v_{m}^{k})\Re(\overline{y}_{m})\Re(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})\!-\!\Im(v_{m}^{k})\Im(\overline{y}_{m})\Im(\bm{a}_{m}^{\mathrm{T}}\hat{\mathbf{x}})+c^{\prime}\bigg$$\bigg\}$, (13)
@@
The third term in Equation (12) is non-convex, but can be relaxed to a convex form via smooth approximation[52]:
@@
where $\eta>0$ is a small constant. A smaller value of $\eta$ provides a closer approximation but may reduce the robustness of the algorithm. Typically, $\eta$ is set to $10^{-6}$. The smoothing technique approximates the original non-smooth, non-convex objective with a differentiable surrogate. While this surrogate remains non-convex, the smoothing facilitates the application of optimization algorithms that require differentiability, such as the IRLS method[53, 52]. Specifically, in the context of IRLS, smoothing allows for the formulation of weight updates and iterative schemes that can converge to stationary points of the original non-convex problem[53].
@@
[…]
where $\mathbf{v}^{k}=[v_{1}^{k},\cdots,v_{M}^{k}]^{\mathrm{T}}\in\mathbb{C}^{M\times 1}$. We have the following theorem regarding the local convergence of $\hat{\mathbf{x}}^{\star}$ in SBRI framework.
Theorem 1.
Let $\{(\hat{\mathbf{x}}^{k},\mathbf{v}^{k})\}_{k=0}^{\infty}$ be the sequence generated by Algorithm 1, and assume the following:
  1. (i)
    The measurement matrix $\mathbf{\mathcal{A}}$ is bounded;
  2. (ii)
    The regularizer $\sum_{i=1}^{N}(|\hat{x}_{i}|^{2}+\eta)^{\alpha/2}$ is finite, lower semicontinuous, and satisfies $0<\alpha\leq 1$ and $\eta>0$;
  3. (iii)
    The sequence $\{\gamma^{k}\}$ remains bounded.
Then the joint Lyapunov function
$\Phi(\hat{\mathbf{x}}^{k},\mathbf{v}^{k}):=\beta\big\|\mathbf{v}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k}\big\|_{2}^{2}+\sum_{i=1}^{M}H(v_{i}^{k})+\frac{\gamma^{k}}{\alpha}\sum_{i=1}^{N}(|\hat{x}_{i}^{k}|^{2}+\eta)^{\alpha/2}$
with $\beta\geq 0$ is non-increasing and bounded. Consequently, the sequence $\{(\hat{\mathbf{x}}^{k},\mathbf{v}^{k})\}$ is bounded and admits at least one convergent subsequence. Moreover, the limit point $(\hat{\mathbf{x}}^{\dagger},\mathbf{v}^{\dagger})$ of any convergent subsequence is a stationary point of the alternating update system.
Corollary 1.
Suppose that the 1-bit quantized observations $\overline{\mathbf{y}}$ are generated from a true signal $\hat{\mathbf{x}}^{\diamondsuit}$ via
$\overline{\mathbf{y}}=\operatorname{csgn}(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{\diamondsuit}+\mathbf{n}),$
where $\hat{\mathbf{x}}^{\diamondsuit}$ is an $s$-sparse vector and $\bf{n}$ is additive noise. Assume that the measurement matrix $\mathbf{\mathcal{A}}$ satisfies the Restricted Isometry Property (RIP) of order $2s$ with constant $\delta_{2s}<1$, and that the smoothing parameter $\eta>0$. Let $\{\hat{\mathbf{x}}^{k}\}$ be the sequence produced by Algorithm 1 with $0<\alpha\leq 1$. Then any convergent subsequence of $\{\hat{\mathbf{x}}^{k}\}$ admits a limit $\hat{\mathbf{x}}^{\dagger}$ that satisfies the recovery bound
$\left\lVert\hat{\mathbf{x}}^{\dagger}-\hat{\mathbf{x}}^{\diamondsuit}\right\rVert_{2}\leq C_{1}+C_{2}\cdot\sigma_{s}(\hat{\mathbf{x}}^{\dagger})_{2},$
where $\sigma_{s}(\hat{\mathbf{x}}^{\dagger})_{2}:=\inf_{\|\mathbf{z}\|_{0}\leq s}\|\hat{\mathbf{x}}^{\dagger}-\mathbf{z}\|_{2}$ denotes the best $s$-sparse approximation error, and $C_{1},C_{2}>0$ are constants depending only on $\delta_{2s}$ and the initial point $\hat{\mathbf{x}}^{0}$.
The proof of Theorem 1 and Corollary 1 is in the Appendix A. Note that Equation (16) represents a classic regularized least squares problem, and $\gamma$ is a regularization parameter. Therefore, we utilize the IRLS method[54], leveraging the estimates obtained from the $k$-th iteration step, to compute $\hat{\mathbf{v}}^{k}$ and $\hat{\mathbf{x}}^{k+1}$ as:
$\mathbf{v}^{k}$ $\!\!=[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{D}^{k}\!\!-\mathrm{I}^{\prime}(\mathbf{D}^{k}))]+j[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{D}^{k}\!\!-\mathrm{I}^{\prime}(\mathbf{D}^{k}))],$ (17)
$\hat{\mathbf{x}}^{k+1}$ $=\left(\mathcal{A}^{\mathrm{H}}\mathcal{A}+\gamma^{k}\bm{\Lambda}(\hat{\mathbf{x}}^{k})\right)^{-1}\mathcal{A}^{\mathrm{H}}\mathbf{v}^{k},$ (18)
where $\mathbf{D}^{k}=\left[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k})\right]+j\left[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{\mathcal{A}}\hat{\mathbf{x}}^{k})\right],$ and
@@
Algorithm 1 SBRI for One-Bit On-Grid DOA
2:Output: $\hat{\mathbf{x}}^{\star}$
4:for $k=0$ to $T_{\max}-1$ do
5:  Compute $\mathbf{v}^{k}$ via (17): $\mathbf{v}^{k}\leftarrow[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{D}^{k}-\mathrm{I}^{\prime}(\mathbf{D}^{k}))]+j[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{D}^{k}-\mathrm{I}^{\prime}(\mathbf{D}^{k}))]$
6:  Update $\hat{\mathbf{x}}^{k+1}$ via (18): $\hat{\mathbf{x}}^{k+1}\leftarrow(\mathcal{A}^{\mathrm{H}}\mathcal{A}+\gamma^{k}\Lambda(\hat{\mathbf{x}}^{k}))^{-1}\mathcal{A}^{\mathrm{H}}\mathbf{v}^{k}$
7:  $\gamma^{k+1}\leftarrow\gamma^{0}\|\hat{\mathbf{x}}^{k+1}\|_{2}$
8:  if convergence($\hat{\mathbf{x}}^{k},\hat{\mathbf{x}}^{k+1},\epsilon_{0}$) then break
9:  end if
10:end for
11:return $\hat{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k+1}$